OpenStudy (anonymous):

Please Help!!! Solve the system of equations 4x-2y=2, -3x+7y=-29.

OpenStudy (3psilon):

4x-2y=2 -3x+7y=-29 Make the x's the same number By multiplying both sides of both equations by the same number 3(4x-2y)=2(3) 4(-3x+7y)=-29(4) Doing that gets you 12x - 6y = 6 -12x + 28y = -116 Combine the like terms X's cancel out 22y = -110 solve for Y then plug in y to either equations to get X